determine the net force necessary to give an object with a mass of 2.0 kg an acceleration of 5.0 m/s2.

Answer by ikleyn(52803)

.
Force  F = m*a = 2.0 kg * 5 m/s^2 = 10 (kg*m)/s^2 = 10 N = 10 newtons.


Newton's second law.